OpenStudy (anonymous):

can someone walk me through this problem? Let f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (anonymous):

f(-1) is just the value of the function f when x = -1.

OpenStudy (anonymous):

Function f is called x^2 - 8x + 5.

OpenStudy (anonymous):

What is the value of f when x = -1?

OpenStudy (anonymous):

ok so i just plug x in and then solve?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok hold on please

OpenStudy (anonymous):

so f = 11?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok ill try again

OpenStudy (anonymous):

Let's see...when x = -1, you get (-1)^2 -8(-1) + 5 1 + 8 + 5 = 14 Agree?

OpenStudy (anonymous):

oh okay i see what happened i forgot the first x was to the second power

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome. Now you can find f(anything).

OpenStudy (anonymous):

how would i solve with the f^-1? would i just turn the number into a negative number? Let f(x) = 2x - 6 Solve f^-1(x) when x = 2.

OpenStudy (anonymous):

@Easyaspi314 ?

OpenStudy (anonymous):

f^(-1) is the inverse of function f.

OpenStudy (anonymous):

To find the inverse of a function f, interchange x with y, and y with x, and then solve for y.

OpenStudy (anonymous):

As an example: What is the inverse of f(x) = 2x + 3? ANSWER: We are given that y = 2x + 3 We interchange x with y and y with x, we get x = 2y + 3 Now, we solve for y, we get y = (x-3)/2 So we say that the inverse of f, f^(-1) = (x-3)/2.